OpenStudy (rohan5599):

sec of theta=9/4, theta in quadrant IV Find tan of theta

OpenStudy (anonymous):

In Q4, cos and sec are positive, everything else is negative. Draw a triangle.

OpenStudy (anonymous):

If sec = (9/4), then cos=(1/(9/4)) = (4/9)

OpenStudy (anonymous):

Solve for the other side to find the tangent, just pretend there is an angle at the left corer.